OpenStudy (anonymous):

whoever answers this gets a medal!!An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. 240 mph 220 mph 200 mph 180 mph

OpenStudy (anonymous):

Headwind distance = 3.5(x-40) Tailwind distance = 2.5(x+40) 3.5(x-40) = 2.5(x+40) 3.5x-140=2.5x+100 x=140+100 x=240 mph speed of plane in still air The answer is A)240

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome jader. anything else u need help with?

OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

please :)